OpenStudy (anonymous):

find the slope of the line 5x-2y=-1

OpenStudy (ybarrap):

5x - 5y = -1 5x + 1 = 5y y = (5x + 1)/5 = x + 1/5 Ans. slope = 1

jimthompson5910 (jim_thompson5910):

5x-2y=-1 -2y=-1-5x -2y=-5x-1 y=-5x/(-2)-1/(-2) y=(5/2)x+1/2 So the slope is 5/2

OpenStudy (anonymous):

the slope is 5/2